OpenStudy (anonymous):

What is the energy of a single photon of visible light with a wavelength of 500.0 nm? 1 nm=10^-9 m

OpenStudy (anonymous):

A. 2.02 * 10^18 J B. 2.84 * 10^14 J C. 3.97 * 10^-19 J D. 4.95 * 10^-19 J

rvc (rvc):

E = h N h : plancks constant N : 500 nm

OpenStudy (michele_laino):

I think that the requested energy E is: \[\large E = \frac{{hc}}{\lambda } = \frac{{6.62 \times {{10}^{ - 34}} \times 3 \times {{10}^8}}}{{500 \times {{10}^{ - 9}}}} = ...joules\]

rvc (rvc):

as usual i mess the formula lol

OpenStudy (anonymous):

ok! we get 3.972 E -19! so choice C is our solution!! thanks!!:)

OpenStudy (michele_laino):

yes! correct!